OpenStudy (anonymous):

Part 1: Solve the inequality −5x + 8 less than or greater to x − 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

I did part one, can someone help me with the last part. -5x+8<=x-4 -5x+8-8<=x-4-8 -5x<=x-12 -5x-x<=x-x-12 -6x<=-12 -6x/-6=>-12/-6 x=>2

OpenStudy (ivettef365):

is it less than or equal to (you wrote less than or greater to)

OpenStudy (ivettef365):

and yes you did it right

OpenStudy (anonymous):

less than equal to sorry

OpenStudy (anonymous):

part 2 basically is put in word how i did it all and what type of circle and which side its shaded in.. im not good at that part .

OpenStudy (anonymous):

@nubeer

OpenStudy (ivettef365):

shading goes to the right

OpenStudy (nubeer):

greater then equal to means.. that a closed circle.

OpenStudy (nubeer):

making a number line .. put numbers.. and the circle.. that would help to start.

OpenStudy (anonymous):

right side shaded right..

OpenStudy (nubeer):

yes

OpenStudy (anonymous):

thanks

OpenStudy (nubeer):

yw :)